OpenStudy (anonymous):

L=c+(e-1)m; solve for e?

OpenStudy (anonymous):

just do like you did, i can check ur answer

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

\[\frac{ c-1 }{ m } +1?\]

OpenStudy (anonymous):

not too sure.

OpenStudy (anonymous):

L=c+(e-1)m L = c + em - m (L-c+m)/m = e L/m - c/m + 1 = e lol your L disappeared again

OpenStudy (anonymous):

lol my bad